# Solving the PTA Data Analysis Problem with a Global Gibbs Scheme

### v1

@@ VI.1 Fiducial model
The most fair comparison of MG to standard detection pipelines would be if the latter were able to infer the cross-power-spectral-density between each pulsar pair. However, as mentioned earlier, the dimensionality of this problem is too high for standard techniques based on Metropolis-Hastings MCMC sampling to confront. Therefore, as a proxy for this in our standard detection pipeline (built with enterprise), we assume a flexible model for the inter-pulsar correlations in which these correlations are the same at all GW frequencies, but are allowed to vary in seven independent angular-separation bins, similar to Figure 1-(d) of Agazie et al. [5]. [^10] The 7 bin locations are chosen to capture the maximum and minimum angular separations, the two zero crossings of the Hellings and Downs curve as well as the position of minimum correlation. We call this a common cross-correlated red noise (CCRN) process; it captures some of the flexibility of MG while still running in a reasonably short amount of computational time. Spectra in this model are treated as power-law across frequencies for the CCRN process and for intrinsic pulsar red noise, with associated power-law amplitude and spectral index parameters. This model is therefore that of Equation 8, with $\Gamma_{IJ}=\Gamma_{IJ}(\bm{c})$ modeled as a piecewise constant function with inferred values in different angular separation ranges.
@@ VI.3.2 Characterizing the cross-spectrum
Figure 4: The distribution of $\log_{10}$ of the optimal estimator of $A_{\text{GWB}}$ found by applying the $\chi^{2}$ statistic of Equation 27 on unique random draws from the posterior of $\Phi_{IJ}$. The lower distribution (purple) is from analyzing simReal, whereas the upper distribution (light orange) is from analyzing simGWB. For both datasets the injected GWB amplitude is $A_{\mathrm{GWB}}=6\times 10^{-15}$.
@@
Figure 5: Left: a density plot of power-spectral density fits to the cross-correlation estimates of MG ($\Phi_{IJ}$ for $I\neq J$) at the first frequency-bin is depicted. Right: a plot of the reconstructed shape of the correlations (median and 1-$\sigma$ level) using fits to $\Phi_{IJ}$ is shown. Both plots are obtained through Levenberg-Marquardt (LM) [45] fits to random draws from simReal’s $\Phi_{1;IJ}$ posterior probability distributions. The fitted model follows Equation 28 with only two parameters: $\log_{10}{\rho_{1}}$ for the power-spectral density and $\tau$ for the shape of the correlations.
@@
We now use the techniques described in §V.2 for fitting inter-pulsar–correlated models to the sampled output of MG. Figure 4 shows the result of applying the $\chi^{2}$ statistic of Equation 27 to the joint posterior on all $\Phi_{IJ}$ elements (with $I\neq J$) that were sampled by MG on both sets of simulated data. To make Figure 4, we randomly draw $2\times 10^{5}$ unique samples from the $\Phi_{IJ}$ MG posterior, noting that these draws respect the measurement covariance between different pulsar pairs. For each draw, the value of $A_{\text{GWB}}$ that minimizes the $\chi^{2}$ statistic of Equation 27 is estimated; hence Figure 4 is a distribution of derived best-fit GWB amplitudes over the MG posterior. We can see that, in these cases, the distributions are highly consistent with the injected amplitude value, which falls almost at the mode of each.
@@
Instead of assuming that the pattern of inter-pulsar correlations conforms to the Hellings and Downs function, we now use the GT model described by Equation 28. The general procedure is outlined in §V.2; in this specific test we use the LM algorithm to find the best-fit power-spectral-density value, $\rho_{1}$ and GT shape parameter, $\tau$, at frequency $f_{1}=1/T_{\text{obs}}$. Figure 5 shows the result of fitting the GT model to $2\times 10^{5}$ unique random draws from the $\Phi_{1;IJ}$ posterior of simReal. The LM algorithm is applied to these draws to find the best-fit model parameters. The left panel of Equation 28 shows the uncertainty distribution on $\log_{10}{\rho_{1}}$, which is consistent with the injected value shown as a dashed vertical line. The best-fit $\tau_{1}$ estimates are used to create a posterior envelope of GT curves, shown as a solid median curve and shaded $1$-$\sigma$ region. The true correlation pattern—i.e., the Hellings and Downs function—is shown as a dashed line that falls well within the shaded envelope of reconstructed inter-pulsar correlation curves.
@@ VII No sampling— approximating the GWB characterization using Inverse-Wishart distribution
Note that the success of the presented approximation method is dependent on the closeness of MG’s Fourier coefficients to the ones obtained from CURN and single-pulsar analyses. For theoretical simulated data sets with strong red noise signal and ideal timing conditions, this is not an issue as Fourier coefficients’ recovery is not much sensitive to the Bayesian models used to infer them. Hence, single-pulsar analyses would yield the same distribution of Fourier coefficients as MG or CURN would. However, for reasons yet to be fully understood, this becomes an issue with realistic data sets: single-pulsar results disagree with PTA-level analyses significantly in both PSD and Fourier coefficient recovery of the red noise of many pulsars.
@@ VIII Conclusion and Future Work
As shown in this work, spectrum fitting techniques using one-dimensional kernel-density-estimators could be used to refit the red noise auto-spectrum samples generated by MG and obtain posterior probability distributions for both common correlated and intrinsic uncorrelated red noise model parameters for each pulsar. Furthermore, we show examples of how one can perform a cross-correlation-only analysis using the cross-spectrum samples generated by MG. The first example involves using a $\chi^{2}$ statistic to find a distribution for the optimal estimator of the amplitude of a powerlaw model with fixed spectral index and fixed correlations at Hellings and Downs values. The powerlaw model whose amplitude is estimated using the $\chi^{2}$ statistic describes the functional form of the cross-correlation power-spectral-density of the GWB, akin to the noise marginalized optimal statistic technique [27, 42]. Additionally, we show how to perform an Levenberg-Marquardt fit (more commonly known as the LM fit) to fit a specific model to the cross-correlation samples of MG; hence, perform inference about the model parameters that characterize the correlations induced by a GWB signal.
@@
The generality of the MG’s noise modeling comes at a price. The price is numerical instability of linear algebra operations that need to be carried to complete every iteration of Gibbs sampling. In this paper, we show that the application of our current implementation of MG on both realistic and theoretical data sets is consistent with the injected signal as well as the results of the standard analysis aimed at characterizing the auto and cross-spectrum of the GWB. Nevertheless, to prepare MG for uses in real data set analyses, we will need to improve on our current implementation MG ensuring its numerical stability at all times during the sampling steps.
@@ VIII.1 Software
The multi-pulsar Gibbs sampling code takes advantage of the functionalities provided by Numpy [48], Scipy, [49], Ceffyl [22], ENTERPRISE [50] and ENTERPRISE-extensions [43]. Python packages matplotlib [51], and chainconsumer [52] have been used for generating the scientific plots in this paper.
@@ Acknowledgements.
We thank our colleagues in NANOGrav for fruitful discussions and feedback during the development of this technique. The work of N.L., W.G.L, X.S., and S.R.T was supported by the NANOGrav NSF Physics Frontier Center awards #2020265 and #1430284. X.S acknowledges the support from the George and Hannah Bolinger Memorial Fund at Oregon State University. S.R.T acknowledges support from NSF AST-2007993, and an NSF CAREER #2146016. N.L is supported by the Vanderbilt Initiative in Data Intensive Astrophysics (VIDA) Fellowship. This work was conducted in part using the resources of the Advanced Computing Center for Research and Education (ACCRE) at Vanderbilt University, Nashville, TN.

### v2

@@ VI.1 Fiducial model
The most fair comparison of MG to standard detection pipelines would be if the latter were able to infer the cross-power-spectral-density between each pulsar pair. However, as mentioned earlier, the dimensionality of this problem is too high for standard techniques based on Metropolis-Hastings MCMC sampling to confront. Therefore, as a proxy for this in our standard detection pipeline (built with the PTA data analysis suite enterprise [51, 48]), we assume a flexible model for the inter-pulsar correlations in which these correlations are the same at all GW frequencies, but are allowed to vary in seven independent angular-separation bins, similar to Figure 1-(d) of Agazie et al. [5]. [^11] The 7 bin locations are chosen to capture the maximum and minimum angular separations, the two zero crossings of the HD curve as well as the position of minimum correlation. We call this a common cross-correlated red noise (CCRN) process; it captures some of the flexibility of MG while still running in a reasonably short amount of computational time. Spectra in this model are treated as power-law across frequencies for the CCRN process and for intrinsic pulsar red noise, with associated power-law amplitude and spectral index parameters. This model is therefore that of Equation 8, with $\Gamma_{IJ}=\Gamma_{IJ}(\bm{c})$ modeled as a piecewise constant function with inferred values in different angular separation ranges.
@@ VI.3.2 Characterizing the cross-spectrum
Figure 4: The distribution of $\log_{10}$ of the optimal estimator of $A_{\text{GWB}}$ found by applying the $\chi^{2}$ statistic of Equation 27 on unique random draws from the posterior of $\Phi_{IJ}$. Note the use of “hat” on $A_{\text{GWB}}$ to emphasize the fact that the distributions are made from best-fit values. The lower distribution (purple) is from analyzing simReal, whereas the upper distribution (light orange) is from analyzing simGWB. For both datasets, the injected GWB amplitude (black vertical dashed line) is $A_{\mathrm{GWB}}=6\times 10^{-15}$. The 1-$\sigma$ level of each distribution is identified with vertical dotted lines matching the color of their corresponding distribution.
@@
Figure 5: Left: a density plot of power-spectral density fits to the cross-correlation estimates of MG ($\Phi_{IJ}$ for $I\neq J$) at the first frequency-bin is depicted. Right: a plot of the reconstructed shape of the correlations (median and 1-$\sigma$ level) using fits to $\Phi_{IJ}$ is shown. Both plots are obtained through LM fits to random draws from simReal’s $\Phi_{1;IJ}$ posterior probability distributions. The fitted model follows Equation 28 with only two parameters: $\log_{10}{\rho_{1}}$ for the power-spectral density and $\tau$ for the shape of the correlations.
@@
We now use the techniques described in §V.2 for fitting inter-pulsar–correlated models to the sampled output of MG. Figure 4 shows the result of applying the $\chi^{2}$ statistic of Equation 27 to the joint posterior on all $\Phi_{IJ}$ elements (with $I\neq J$) that were sampled by MG on both sets of simulated data. To make Figure 4, we randomly draw $2\times 10^{5}$ unique samples from the $\Phi_{IJ}$ MG posterior, noting that these draws respect the measurement covariance between different pulsar pairs. For each draw, the value of $A_{\text{GWB}}$ that minimizes the $\chi^{2}$ statistic of Equation 27 is estimated (denoted by $\hat{A}_{\text{GWB}}$); hence Figure 4 is a distribution of derived best-fit GWB amplitudes over the MG posterior. We can see that, in these cases, the distributions are consistent with the injected amplitude value to within the 1-$\sigma$ interval of the distributions.
@@
Instead of assuming that the pattern of inter-pulsar correlations conforms to the HD curve, we now use the GT model described by Equation 28. The general procedure is outlined in §V.2; in this specific test we use the LM algorithm to find the best-fit power-spectral-density value, $\rho_{1}$ and GT shape parameter, $\tau$, at frequency $f_{1}=1/T_{\text{obs}}$. Figure 5 shows the result of fitting the GT model to $2\times 10^{5}$ unique random draws from the $\Phi_{1;IJ}$ posterior of simReal. The LM algorithm is applied to these draws to find the best-fit model parameters. The left panel of Equation 28 shows the uncertainty distribution on $\log_{10}{\rho_{1}}$, which is consistent with the injected value shown as a dashed vertical line. The best-fit $\tau_{1}$ estimates are used to create a posterior envelope of GT curves, shown as a solid median curve and shaded $1$-$\sigma$ region. The true correlation pattern—i.e., the HD curve—is shown as a dashed line that falls in the shaded envelope of the reconstructed inter-pulsar correlation curves.
@@ VII No sampling— approximating the GWB characterization using Inverse-Wishart distribution
Note that the success of the presented approximation method is dependent on the closeness of MG’s Fourier coefficients to the ones obtained from CURN and single-pulsar analyses as mentioned previously. For theoretical simulated data sets with strong red noise signal and ideal timing conditions, this is not an issue as Fourier coefficients’ recovery is not much sensitive to the Bayesian models used to infer them. Hence, single-pulsar analyses would yield the same distribution of Fourier coefficients as MG or CURN would. However, for reasons yet to be fully understood, this becomes an issue with realistic data sets: single-pulsar results disagree with PTA-level analyses significantly in both PSD and Fourier coefficient recovery of the red noise of many pulsars.
@@ VIII Conclusion and Future Work
As shown in this work, spectrum fitting techniques using one-dimensional kernel-density-estimators could be used to refit the red noise auto-spectrum samples generated by MG and obtain posterior probability distributions for both common correlated and intrinsic uncorrelated red noise model parameters for each pulsar. Furthermore, we show examples of how one can perform a cross-correlation-only analysis using the cross-spectrum samples generated by MG. The first example involves using a $\chi^{2}$ statistic to find a distribution for the optimal estimator of the amplitude of a powerlaw model with fixed spectral index and fixed correlations at HD values. The powerlaw model whose amplitude is estimated using the $\chi^{2}$ statistic describes the functional form of the cross-correlation power-spectral-density of the GWB, akin to the noise marginalized optimal statistic technique [29, 47]. Additionally, we show how to perform an LM fit to fit a specific model to the cross-correlation samples of MG; hence, perform inference about the model parameters that characterize the correlations induced by a GWB signal.
@@
The generality of the MG’s noise modeling comes at a price. The price is numerical instability of linear algebra operations that need to be carried to complete every iteration of Gibbs sampling. In this paper, we show that the application of our current implementation of MG on both realistic and theoretical data sets is consistent with the injected signal as well as the results of the standard analysis aimed at characterizing the auto and cross-spectrum of the GWB. Nevertheless, to prepare MG for analyzing real data set, we will need to improve on our current implementation of MG ensuring its numerical stability and scalability at all times during the sampling steps. The improvements are needed as reaching adequate levels of convergence for the Fourier coefficients becomes more and more difficult with increasing number of frequency-bins and pulsars with mostly white noise dominated data. The scalability of MG to include more pulsars and frequency-bins may be resolvable by using modern GPU-based array-oriented numerical computation software such as JAX [54] to reduce the computational cost of steps shown in Figure 1. Additionally, we speculate that the most effective solution to resolve the numerical problems of MG is possible through constrained draws from the multivariate normal distribution of Equation 13 to prevent the Gibbs sampler from getting stuck at regions of parameter space that correspond to unreasonably weak ($<10^{-8}$ seconds) Fourier coefficient estimate.
@@ VIII.1 Software
The multi-pulsar Gibbs sampling code [55] takes advantage of the functionalities provided by Numpy [56], Scipy, [57], Ceffyl [24], enterprise [51] and enterprise-extensions [48]. Python packages matplotlib [58], and chainconsumer [59] have been used for generating the scientific plots in this paper. The simulated data sets can be found in Laal [55].
@@ Acknowledgements.
We thank the anonymous referee for their helpful feedback which improved the quality of this work. We thank Bernard Whiting for pointing out plotting errors in some of the figures featured in this work. We thank our colleagues in NANOGrav for fruitful discussions and feedback during the development of this technique. The work of N.L., W.G.L, X.S., and S.R.T was supported by the NANOGrav NSF Physics Frontier Center awards #2020265 and #1430284. X.S acknowledges the support from the George and Hannah Bolinger Memorial Fund at Oregon State University. S.R.T acknowledges support from NSF AST-2007993, and an NSF CAREER #2146016. N.L is supported by the Vanderbilt Initiative in Data Intensive Astrophysics (VIDA) Fellowship. This work was conducted in part using the resources of the Advanced Computing Center for Research and Education (ACCRE) at Vanderbilt University, Nashville, TN.